Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Difficult Math Problem #85 - Probability

Expert replies
by 800guy » Wed Jan 10, 2007 9:06 am
Brad flips a two-sided coin 8 times. What is the probability that he gets tails on at least 7 of the 8 flips?

A) 1/32

B) 1/16

C) 1/8

D) 7/8

E) none of the above


OA coming after a few people try to answer. Source: Difficult Math Problems doc on this forum
Join the discussion
Source: — Problem Solving |

by Argen » Wed Jan 10, 2007 10:49 am
Using the bernoulli distribution, lets add p(getting 7 tails) + p(getting 8 tails):

(8C7)*(0.5^7)*(0.5) + (8C8)*(0.5^8) = 8/(2^8) + 1/(2^8) = 9/256

It's also the same as the probability of getting at most 1 head. We calculate p(at most 1 head) = p(0 heads) + p(1 head):

(8C0)*(0.5^0)*(0.5^8) + (8C1)*(0.5^1)*(0.5^7) = 1/(2^8) + 8/(2^8) = 9/256
Join the discussion

OA

by 800guy » Fri Jan 12, 2007 12:31 am
OA

Number of ways 7 tails can turn up = 8C7
the probability of those is 1/2 each
Since the question asks for at least 7, we need to find the prob of all 8 tails - the number of ways is 8C8 = 1
Add the two probabilities
8C7*(1/2)^8 = 8/2^8 -- for getting 7
Prob of getting 8 tails = 1/2^8
Total prob = 8/2^8+1/2^8 = 9/2^8

Ans is E.
Join the discussion

i think i am being asked for probability of 7 tails & probability of getting 8 tails in 8 throws
total possible outcomes=2(possible outcomes on any throw)*8(total throws)
=16

probability that he gets tails on at least 7 of the 8 flips=7/16*8/16=7/32
but this is wrong answer. can someone suggest how to calculate the right answer without combinatorics or suggest how to use Bernoulli distribution (i do not understand whether it is used to combine probabilities or it is used after we calculate total possible and favorable events)
Join the discussion

by stop@800 » Sun Sep 21, 2008 12:29 pm
Ashish, others,

Answer without formulas:

Totals 8 flips: possible arrangements = 2^8

we need 7 tails and 8 tails
in 8 ways we can have 7 tails [as 1 h can be at 8 places]
and
in 1 way we can have 8 tails

so favorable outcomes = 8+1 = 9
Total outcomes = 2^8 = 256
Join the discussion

pl help

by stop@800 » Sun Sep 21, 2008 12:36 pm
Where can I find
Source: Difficult Math Problems doc on this forum

Thanks
Join the discussion